What are facts about the number 46?
In chemistry, 46 is the atomic number of Palladium.
In human biology, there are 46 human chromosomes. There are 22 pairs of autosomes and one pair of sex chromosomes.
What are the Probable Sources to E. coli?
the most probable sources to e-coli disease is eating contaminated food because that is where most cases originate from. that and maybe your mom gave you some bad cooking.
Assuming that the man who has normal vision is homozygous for normal vision, the couple's daughter will either be homozygous for normal vision or heterozygous (normal vision but carrier for color blindness) for normal vision. In light of this, the couple's daughter will not be color blind.
The probability of drawing a seven from a standard deck of 52 cards is?
There are 4 sevens so probability of drawing a seven is 4/52 or 1/13 or 0.0769.
What is the probability of Drawing 3 aces from a deck of cards?
The probability of drawing three aces from a deck of cards is 1 in 5525.
The probability of the first ace is 4 in 52, or 1 in 13. The second ace is 3 in 51, or 1 in 17. The third ace is 2 in 50, or 1 in 25. Multiply these three probabilities together and you get 1 in 5525.
What are the odds of a child having the same birthday as the mother?
Believe it or not it happens very often for example three generations of women in my family have the same birthday i am the third generation I've seen it a lot more too so the odds are high.
I agree, I am born on my mother's birthday and in school I had several friends who were also born on their mother's birthdays.
What is the probability of rolling a sum of 15 on two fair dice?
Zero is the probability of rolling a sum of 15 on two fair dice; the maximum value is 12.
What can you infer about why the probability of an earthquake is so high in the Parkfield area?
You can infer that it is on or near a fault line. You can also infer that a earthquake is very likely to occur there.
What is the definition for Probability?
Probability is described as the likelihood of a particular event happening. For example, say you are betting on a horse race, each horse has a particular probability of winning.
The likelihood of an event occuring
The proportion of times an event occurs over a large number of trials
A ratio of successful outcomes to total possible outcomes
For a random event, the proportion of times an event occurs over a large nuber of trials
How does Mugenda and Mugenda define sample size?
According to research made at multiple websites; Mugenda and Mugenda, which is a company that focuses on research methods for various areas, define sample size by using statistics data and probability.
What is the probability of dealing 4 diamonds in a row out of a deck of cards?
(13/52)*(12/51)*(11/50)*(10/49) = 17160/6497400
The probability is 0.3692
What 4 kings do a deck of cards represent?
The king of spades represents David,King of Israel. The king of clubs is said to represent Alexander the Great. The king of hearts represents the French king Charlemogne, and the king of diamonds represents Roman Ceaser Augustus.
When a deck of cards slips out of your hands and fall to the ground gravity does work on it?
True - apex =)
It is 2/11.
What is the probability of drawing and ace or a club?
aces in card=4
P(Aces)=4/52
clubs are= 13
P(Club) = 13/52
P(An Ace or a Club)= P(Ace). P (club)= 4/52*13/52=1/52
You cannot be sure.
Not sure how the faces of a coin are similar to genes. There are two and you will get one from each parent. One or the other will dominate. However, while heads and tails of a coin are equally "powerful", some genes are far more powerful (recessive) than others.
2/5 or 40%.